OpenStudy (anonymous):

what is square root of 6 + sqroot of 6

OpenStudy (slaaibak):

2sqrt6

OpenStudy (anonymous):

\[\sqrt{6}+\sqrt{6}\]

OpenStudy (anonymous):

is it\[\sqrt{12}\]

OpenStudy (pokemon23):

2.906800603

OpenStudy (anonymous):

thats it pokemon

OpenStudy (anonymous):

It is not \(\sqrt{12}\). It can be written as either \(2\sqrt{6}\), \(\sqrt{24}\), or as 2.906....

OpenStudy (slaaibak):

If sqrt6 = x what is x+x? Apply that logic to this

OpenStudy (pokemon23):

TI-84 Plus Silver Edition never lies ....

OpenStudy (anonymous):

Scratch that 2.906..., Lillia has the correct decimal version.

OpenStudy (anonymous):

ok so the answers are a. \[2\sqrt{6}\] b. 6 c. \[\sqrt{12}\] d. 36

OpenStudy (anonymous):

add or subtract as indicated

OpenStudy (anonymous):

a.

OpenStudy (pokemon23):

a

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Slaaibak said it rather well: "If sqrt6 = x what is x+x? Apply that logic to this"

OpenStudy (anonymous):

ok ive been doing math for 7 hrs now My brain is fried.